# Non-abelian Hodge correspondence over singular Kähler spaces

### v1

@@ Abstract.
The proof relies on two key ingredients: first, we establish an equivalence over the regular loci—via harmonic bundles—between polystable Higgs bundles with vanishing orbifold Chern numbers and semi-simple flat bundles; second, we prove a descent theorem for semistable Higgs bundles with vanishing Chern classes along resolutions of singularities.
@@ 1. Introduction
$\left\{\begin{subarray}{c}\text{polystable Higgs bundles}\\ \text{with vanishing Chern numbers}\end{subarray}\right\}$$\left\{\text{harmonic bundles}\right\}$$\left\{\begin{subarray}{c}\text{semi-simple linear representations}\\ \text{from }\pi_{1}(X)\text{ to }\mathrm{GL}(r,\mathbb{C})\end{subarray}\right\}$$\left\{\text{semi-simple flat bundles}\right\}$}Hitchin-SimpsonNon-abelian Hodge correspondenceCorlette-DonaldsonRiemann-Hilbert
In general, Simpson [99] established an equivalence between the category of semistable Higgs sheaves with vanishing Chern numbers and the category of linear representations of the fundamental group. This equivalence also holds for compact Kähler manifolds [86, 23] and has been significantly generalised to other settings, such as quasi-Kähler manifolds [98, 61, 83, 84, 85] and varieties over fields of positive characteristic [26, 87, 1, 69, 68, 96], we refer to [74, 52] for an overview of its development.
@@
The main results of this paper establish the non-abelian Hodge correspondence for compact Kähler spaces with Kawamata log terminal (klt) singularities, as well as over their regular loci. This work is motivated by two recent lines of progress: first, the successful construction of the correspondence for projective klt varieties [38, 39]; and second, the recent characterization of stable reflexive sheaves satisfying the orbifold Bogomolov–Gieseker inequality on compact Kähler klt spaces [90, 28] (see also [41, 42]). A further motivation comes from applications to the uniformization of Kähler klt spaces, extending analogous developments in the projective setting (cf. [39, 14, 92, 55] and references therein).
@@ 1.1. Main results
Notation 1.1.
Let $\mathcal{E}_{X_{\text{reg}}}$ be a coherent sheaf defined only on the regular locus of $X$. We denote by $\mathcal{E}_{X}$ its trivial extension to the whole space. Conversely, if $\mathcal{E}_{X}$ is a sheaf on $X$, we write $\mathcal{E}_{X_{\text{reg}}}$ for its restriction to the regular locus. For a morphism $f:Y\to X$ and a coherent sheaf $\mathcal{E}_{X}$ on $X$, we write $f^{[*]}\mathcal{E}_{X}:=(f^{*}\mathcal{E}_{X})^{\vee\vee}$ for the reflexive pull-back (i.e., the double dual of the pull-back).
@@
Let $X$ be a compact Kähler klt space of dimension $n$, we define
  • $\mathrm{Higgs}_{X}$: locally free Higgs sheaves $(\mathcal{E}_{X},\theta_{X})$ of rank $r$ on $X$ such that $(\mathcal{E}_{X},\theta_{X})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable with vanishing Chern classes
@@
Analogously, we can define the categories $\mathrm{pHiggs}_{X}$ and $\mathrm{sLSys}_{X}$ for polystable locally free Higgs sheaves satisfying (1.1) and semi-simple local systems.
The first main result of this paper is establishing the non-abelian Hodge correspondence over compact Kähler klt spaces.
@@
$\mu_{X}:\mathrm{Higgs}_{X}\longrightarrow\mathrm{LSys}_{X}$
@@
commutes. Here $\mu_{\widehat{X}}$ denotes the classical non-abelian Hodge correspondence for the compact Kähler manifold $\widehat{X}$.
@@
We refer to Remark 2.12 for the definition of the pullback $\pi^{*}(\mathcal{E}_{X},\theta_{X})$. Moreover, the restriction of $\mu_{X}$ to $\mathrm{pHiggs}_{X}$ yields a one-to-one correspondence from $\mathrm{pHiggs}_{X}$ to $\mathrm{sLSys}_{X}$ that is likewise compatible with any resolution of singularities.
It is necessary to consider a distinct correspondence $\mu_{X_{\mathrm{reg}}}$ over the regular locus of a compact Kähler klt space $X$, because the induced homomorphism $i_{*}:\pi_{1}(X_{\mathrm{reg}})\to\pi_{1}(X)$ is surjective (cf. [65, Proposition 2.10]) but generally not injective.
@@
Let $X$ be a compact Kähler klt space $X$ of dimension $n$, we define
  • $\mathrm{Higgs}_{X_{\mathrm{reg}}}$: reflexive Higgs sheaves $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ of rank $r$ on $X_{\mathrm{reg}}$ such that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable with vanishing orbifold Chern classes
@@
  • $\mathrm{LSys}_{X_{\mathrm{reg}}}$: local systems of rank $r$ on $X$.
@@
Our definitions employ the orbifold Chern classes, denoted $\widehat{\mathrm{ch}}_{1}$ and $\widehat{\mathrm{ch}}_{2}$, which coincide with $\mathbb{Q}$-Chern classes upon intersection with ample divisors. In particular, the definition of $\mathrm{Higgs}_{X_{\mathrm{reg}}}$ coincides with [39] when $\omega_{0}],\cdots,[\omega_{n-2}]$ are ample. We refer to [90, Section 2] for an introduction of the precise definitions of orbifold Chern classes.
The second main result of this paper is establishing the non-abelian Hodge correspondence over the regular loci of compact Kähler klt spaces.
@@
such that the underlying $C^{\infty}$-bundle of $\mathcal{E}_{X_{\mathrm{reg}}}$ admits a flat connection induced by $\rho_{X_{\mathrm{reg}}}$. Moreover, this correspondence is compatible with the correspondence from Theorem 1.3 in the following sense: for any maximally quasi-étale cover $\gamma:Y\to X$, the reflexive pullback $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ extends to a locally free Higgs sheaf $(\mathcal{E}_{Y},\theta_{Y})$ belonging to $\mathrm{Higgs}_{Y}$, and the diagram
@@
commutes.
@@
Here, $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ denotes the reflexive pullback (cf. Lemma 2.10).
  1. (i)
    As in [38, 39], maximally quasi-étale covers take a central role in the construction of $\mu_{X}$ and $\mu_{X_{\mathrm{reg}}}$. We refer to Section 4.1 for its definitions.
  2. (ii)
    The restriction of $\mu_{X_{\mathrm{reg}}}$ to $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$ gives the bijection from $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$ to $\mathrm{sLSys}_{X_{\mathrm{reg}}}$ that is compatible with every maximally quasi-étale cover.
Together, Theorems 1.3 and 1.7 provide the analytic counterpart of the two correspondences established for projective klt varieties in [38] and [39], respectively.
@@ 1.2. Applications to quasi-uniformization problems
We now present several applications of our main results. First, we give a numerical characterization of singular quotients of complex tori, established via different methods in [14, 42] for a single Kähler form (cf. [75, 35, 13] for previous advances). The smooth version of this result originates in Yau’s solution of the Calabi conjecture [109].
@@
with equality if and only if $X$ is a singular quotient of a torus $\mathbb{T}^{n}$; i.e., there exists a finite group $G\subset\operatorname{Aut}(\mathbb{T}^{n})$ acting freely in codimension one such that $X\cong\mathbb{T}^{n}/G$.
When $K_{X}$ is nef, inequality (1.5)—known as Miyaoka’s inequality, originating from [82]—continues to hold in the Kähler klt setting (cf. [53, Theorem 1.5]). Its equality case has been characterized for compact Kähler manifolds [88, 54, 79] and for projective klt varieties [55]. Extending these results to compact Kähler klt spaces requires a characterization of the semistable reflexive Higgs sheaves that satisfy the orbifold Bogomolov–Gieseker equality [53, Question 7.5], which has been proved for stable reflexive Higgs sheaves in [28] (see also [41, 42]). We now provide such a characterization based on Theorems 1.3 and 1.7, combined with the preservation of semistability under tensor products proved in [110].
@@
Let $X$ be a compact Kähler klt space and let $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover. Suppose $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is a reflexive Higgs sheaf of rank $r$ on the regular locus, semistable with respect to some Kähler forms $\omega_{0},\dots,\omega_{n-2}$. Then the following orbifold Bogomolov–Gieseker inequality holds:
@@
As another key application, we prove a quasi‑uniformization theorem for projective klt varieties with big canonical divisor. The Miyaoka–Yau inequality, originating in [81, 82], is a cornerstone in the geometry of surfaces of general type; see the introductions of [37, 43] for an overview. For a minimal projective klt variety $X$ of general type, the authors of [37] established the orbifold Miyaoka-Yau inequality
@@
This has been extended to minimal projective dlt pairs [43] and, very recently, to minimal Kähler klt spaces [110] (see also [53]). Moreover, the equality case of (1.7) yields a quasi‑uniformization description [39]: the canonical model of $X$ is a singular quotient of the unit ball.
Making use of the framework developed in this paper, we now extend this picture to varieties whose canonical divisor is big, but not necessarily nef.
@@
If equality holds, then the canonical model $X_{\mathrm{can}}$ of $X$ is a singular quotient of a projective manifold $Y$ that is covered by the unit ball $\mathbb{B}^{n}$.
The inequality (1.8) was established in the recent work [53]—a result that is new even in the smooth case. Here $\langle K_{X}^{n-2}\rangle$ denotes the non-pluripolar product; the precise definition of the intersection number appearing in (1.8) is given in [53, Section 4.1]. Together with the very recent preprint [58]—which proves that for a $K$-stable projective klt variety with big anti-canonical divisor, the Miyaoka-Yau type equality forces its canonical model to be a singular quotient of $\mathbb{P}^{n}$—these results confirm that the inequality (1.8) developed in [53] provides an effective criterion for projective klt varieties with big canonical or anti-canonical divisors.
@@ 1.3. Overview and main technical components
Let $X$ be a compact Kähler klt space and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ a reflexive Higgs sheaf on $X_{\mathrm{reg}}$.
@@
    $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{Higgs}_{X_{\mathrm{reg}}}$ (resp. $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$) if and only if for every maximally quasi-étale cover $\gamma:Y\rightarrow X$, $\gamma^{[*]}\mathcal{E}_{X}$ is locally free and $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ extends to a locally free Higgs sheaf $(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{Higgs}_{Y}$ (resp. $\mathrm{pHiggs}_{Y}$). Moreover, $(\mathcal{E}_{Y},\theta_{Y})$ is an extension of locally free Higgs sheaves belonging to $\mathrm{pHiggs}_{Y}$.
@@
    If $\mathcal{E}_{X}$ is locally free, then $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$ (resp. $\mathrm{pHiggs}_{X}$) if and only if $\pi^{*}(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{\widehat{X}}$ (resp. $\mathrm{pHiggs}_{\widehat{X}}$) for every resolution of singularities $\pi:\widehat{X}\rightarrow X$.
The main task is to prove that for any Jordan–Hölder filtration $\{\mathcal{E}_{i}\}_{i=0}^{l}$ of $\gamma^{[*]}(\mathcal{E}_{Y_{\mathrm{reg}}},\theta_{Y_{\mathrm{reg}}})$, each quotient sheaf $\mathcal{E}_{i}/\mathcal{E}_{i-1}$ is locally free and has vanishing Chern classes. Following the spirits of [38, Proposition 7.10], [48, Theorem 1.2] and [55, Proposition 3.10], the problem in our context reduces to proving the following statement: if $(\mathcal{Q}_{Y_{\mathrm{reg}}},\theta_{\mathcal{Q}_{Y_{\mathrm{reg}}}})\in\operatorname{pHiggs}_{Y_{\mathrm{reg}}}$ is stable, then its reflexive extension $\mathcal{Q}_{Y}$ is locally free and all its Chern classes vanish. In the projective case [38, Proposition 7.10], this is handled by cutting down with hypersurface sections to reduce to surfaces, where reflexive sheaves are automatically locally free—a technique unavailable for general Kähler spaces. The central difficulty in the Kähler setting is therefore precisely to establish the local freeness of $\mathcal{Q}_{X}$ in higher dimensions.
In Sections 3 and 4, we concentrate on proving the following characterization of polystable reflexive Higgs sheaves satisfying (1.3). This result not only yields the desired local freeness of $\mathcal{Q}_{Y}$ but simultaneously constructs the correspondence $\mu_{X_{\mathrm{reg}}}:\mathrm{pHiggs}_{X_{\mathrm{reg}}}\to\mathrm{sRep}_{X_{\mathrm{reg}}}$.
@@
    A flat bundle $(E_{X_{\mathrm{reg}}},D_{X_{\mathrm{reg}}})$ is semi-simple if and only if it admits a pluri-harmonic metric $H_{X_{\mathrm{reg}}}$, which is unique up to a positive constant on each simple component.
@@
    A reflexive Higgs sheaf $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ belongs to $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$ if and only if $\mathcal{E}_{X_{\mathrm{reg}}}$ is locally free and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ arises from a harmonic bundle $(E_{X_{\mathrm{reg}}},D_{X_{\mathrm{reg}}},H_{X_{\mathrm{reg}}})$; here $H_{X_{\mathrm{reg}}}$ is unique up to obvious ambiguity. This is also equivalent to the one of the following three statements:
@@
    In Section 3, we prove the existence of a pluriharmonic metric for reflexive Higgs sheaves, which had been proceeded in [28] for the case where $\theta_{X_{\mathrm{reg}}}$ is trivial (see also [41, 42]). Our proof adapts the Higgs version [73] of Bando-Siu’s argument [3] to study the Hermitian-Yang-Mills flow $\{H_{\mathrm{orb}}(t)\}$ on a partial orbifold modification constructed in [67]. Combining this with (1.3), we conclude that the limiting Hermitian–Einstein metric is pluri-harmonic. The overall strategy are detailed in Section 3.1.
@@
    In the first part of Section 4.2, we prove part (1) of the theorem. The existence of pluri-harmonic metrics on semi-simple flat bundles over the regular locus follows from the work of Corlette and Donaldson [17, 25] via maximally quasi‑étale covers; the remaining properties are obtained through the maximum principle, which is well‑adapted to our setting thanks to the extension theorem of Grauert and Remmert [32].
@@
    From Step 1 and Theorem 1.13 (1), we conclude that $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ admits a pluri-harmonic metric induced by a semi-simple local system $E_{Y_{\mathrm{reg}}}$. Since $Y$ is maximally quasi-étale, $E_{Y_{\mathrm{reg}}}$ extends to a semi-simple local system $E_{Y}$. To verify that $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ extends to a locally free Higgs sheaf on $Y$, it suffices to establish a descent result for Higgs bundles arising from harmonic bundles (Proposition 4.11), which we prove using the period map constructed in [19]. Details are given in Sections 4.2 and 4.3.

### v2

@@ Abstract.
The proof relies on two key ingredients: first, we establish an equivalence over the regular loci–via harmonic bundles–between polystable Higgs bundles with vanishing orbifold Chern numbers and semisimple flat bundles; second, we prove a descent (resp. ascent) result for semistable locally free Higgs sheaves with vanishing Chern classes along resolutions of singularities.
@@ 1. Introduction
$\left\{\begin{subarray}{c}\text{polystable Higgs bundles}\\ \text{with vanishing Chern numbers}\end{subarray}\right\}$$\left\{\text{harmonic bundles}\right\}$$\left\{\begin{subarray}{c}\text{semisimple linear representations}\\ \text{from }\pi_{1}(X)\text{ to }\mathrm{GL}(r,\mathbb{C})\end{subarray}\right\}$$\left\{\text{semisimple flat bundles}\right\}$}Hitchin-SimpsonNon-abelian Hodge correspondenceCorlette-DonaldsonRiemann-Hilbert
In general, Simpson [102] established an equivalence between the category of semistable Higgs sheaves with vanishing Chern numbers and the category of linear representations of the fundamental group over a projective manifold $X$. This equivalence also holds for compact Kähler manifolds [88, 23] and has been significantly generalised to other settings, such as quasi-Kähler manifolds [101, 62, 84, 85, 86] and varieties over fields of positive characteristic [26, 89, 1, 70, 69, 99]. We refer to [75, 53] for an overview of its development.
@@
The main results of this paper establish the non-abelian Hodge correspondence for compact Kähler spaces with Kawamata log terminal (klt) singularities, as well as over their regular loci. This work is motivated by two recent lines of progress: first, the successful construction of the correspondence for projective klt varieties [38, 39]; and second, the recent characterization of stable reflexive sheaves satisfying the orbifold Bogomolov-Gieseker inequality on compact Kähler klt spaces [92, 28] (see also [41, 42]). A further motivation comes from applications to the uniformization of Kähler klt spaces, extending analogous developments in the projective setting (cf. [39, 14, 94, 56, 87] and references therein).
@@ 1.1. Main results
Notation 1.1 (Global notations).
For a normal analytic space $X$, we denote by $X_{\mathrm{reg}}$ its regular locus. Let $\mathcal{E}_{X_{\text{reg}}}$ be a coherent sheaf defined only on $X_{\mathrm{reg}}$. We denote by $\mathcal{E}_{X}$ its trivial extension to the whole space. Conversely, if $\mathcal{E}_{X}$ is a sheaf over $X$, we write $\mathcal{E}_{X_{\text{reg}}}$ for its restriction to the regular locus. For a morphism $f:Y\to X$ and a coherent sheaf $\mathcal{E}_{X}$ over $X$, we write $f^{[*]}\mathcal{E}_{X}:=(f^{*}\mathcal{E}_{X})^{\vee\vee}$ for the reflexive pullback (i.e., the double dual of the pullback).
@@
Let $X$ be a compact Kähler klt space of dimension $n$. We define
  • $\mathrm{Higgs}_{X}$: locally free Higgs sheaves $(\mathcal{E}_{X},\theta_{X})$ of rank $r$ over $X$ such that $(\mathcal{E}_{X},\theta_{X})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable with vanishing Chern classes
@@
Analogously, we can define the categories $\mathrm{pHiggs}_{X}$ and $\mathrm{sLSys}_{X}$ for polystable locally free Higgs sheaves satisfying (1.1) and for semisimple local systems, respectively.
The first main result of this paper establishes the non-abelian Hodge correspondence for compact Kähler klt spaces.
@@
$\mu_{X}:\mathrm{Higgs}_{X}\longrightarrow\mathrm{LSys}_{X},$
@@
commutes. Here $\mu_{\widehat{X}}$ denotes the classical non-abelian Hodge correspondence for the compact Kähler manifold $\widehat{X}$. Moreover, the restriction of $\mu_{X}$ to $\mathrm{pHiggs}_{X}$ yields a one-to-one correspondence between $\mathrm{pHiggs}_{X}$ and $\mathrm{sLSys}_{X}$, which is likewise compatible with any resolution of singularities.
@@
We refer to Remark 2.12 for the definition of the pullback $\pi^{*}(\mathcal{E}_{X},\theta_{X})$.
It is necessary to consider a separate correspondence $\mu_{X_{\mathrm{reg}}}$ on the regular locus $X_{\mathrm{reg}}$ of a compact Kähler klt space $X$, because the homomorphism $i_{*}:\pi_{1}(X_{\mathrm{reg}})\to\pi_{1}(X)$ induced by the inclusion map $i_{*}:X_{\mathrm{reg}}\rightarrow X$ is surjective (cf. [66, Proposition 2.10]) but generally not injective.
@@
Let $X$ be a compact Kähler klt space $X$ of dimension $n$. We define
  • $\mathrm{Higgs}_{X_{\mathrm{reg}}}$: reflexive Higgs sheaves $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ of rank $r$ over $X_{\mathrm{reg}}$ such that $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable with vanishing orbifold Chern classes
@@
  • $\mathrm{LSys}_{X_{\mathrm{reg}}}$: local systems of rank $r$ on $X_{\mathrm{reg}}$.
@@
Our definitions employ the orbifold Chern classes, denoted $\widehat{\mathrm{ch}}_{1}$ and $\widehat{\mathrm{ch}}_{2}$, which coincide with $\mathbb{Q}$-Chern classes upon intersection with ample divisors. In particular, the definition of $\mathrm{Higgs}_{X_{\mathrm{reg}}}$ coincides with that in [39] when $\omega_{0}],\cdots,[\omega_{n-2}]$ are ample. We refer to [92, Section 2] for the definitions of orbifold Chern classes (see also [42, Section 5]).
The second main result of this paper establishes the non-abelian Hodge correspondence for the regular loci of compact Kähler klt spaces.
@@
such that the underlying $C^{\infty}$-bundle of $\mathcal{E}_{X_{\mathrm{reg}}}$ admits a flat connection induced by $\rho_{X_{\mathrm{reg}}}$. Moreover, this correspondence is compatible with the correspondence $\mu_{X}$ in the following sense: for any maximally quasi-étale cover $\gamma:Y\to X$, the reflexive pullback $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ extends to a locally free Higgs sheaf $(\mathcal{E}_{Y},\theta_{Y})$ belonging to $\mathrm{Higgs}_{Y}$, and the diagram
@@
commutes. The restriction of $\mu_{X_{\mathrm{reg}}}$ to $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$ gives the bijection from $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$ to $\mathrm{sLSys}_{X_{\mathrm{reg}}}$ that is compatible with every maximally quasi-étale cover; see Theorem 1.13 for a precise statement.
@@
$\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ denotes the reflexive pullback (cf. Lemma 2.10) and we refer to Section 5.3 for the definition of the map $\gamma^{*}:\mathrm{Higgs}_{X_{\mathrm{reg}}}\rightarrow\mathrm{Higgs}_{Y}$ (resp. $\gamma^{*}:\mathrm{LSys}_{X_{\mathrm{reg}}}\rightarrow\mathrm{LSys}_{Y}$). As in [38, 39], maximally quasi-étale covers take a central role in the construction of $\mu_{X}$ and $\mu_{X_{\mathrm{reg}}}$. We refer to Section 4.1 for their definition.
Together, Theorems 1.3 and 1.7 provide the analytic counterparts of the two correspondences established for projective klt varieties in [38] and [39], respectively.
@@ 1.2. Applications to quasi-uniformization problems
We now present several applications of our main results. First, we give a numerical characterization of singular quotients of complex tori, already established in [14, 42] when $\omega_{1}]=\cdots=[\omega_{n-2}]$ (see [76, 35, 13] for earlier developments). The smooth version of this result originates in Yau’s solution of the Calabi conjecture [112].
@@
with equality if and only if $X$ is a singular quotient of a complex torus $\mathbb{T}^{n}$; i.e., there exists a finite group $G\subset\operatorname{Aut}(\mathbb{T}^{n})$ acting freely in codimension one such that $X\cong\mathbb{T}^{n}/G$.
When $K_{X}$ is nef, inequality (1.5)—known as Miyaoka’s inequality, originating from [83]—continues to hold in the Kähler klt setting (cf. [54, Theorem 1.5]). Its equality case has been characterized for compact Kähler manifolds [90, 55, 80] and for projective klt varieties [56]. Extending these results to compact Kähler klt spaces requires a characterization of the semistable reflexive Higgs sheaves that satisfy the orbifold Bogomolov-Gieseker equality [54, Question 7.5], which has been proved for stable reflexive sheaves in [28] (see also [41, 42]). We now establish such a characterization based on Theorems 1.3 and 1.7, combined with the preservation of semistability under tensor products proved in [113].
@@
Let $X$ be a compact Kähler klt space and let $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover. Suppose $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ is a reflexive Higgs sheaf of rank $r$ on the regular locus, semistable with respect to some Kähler forms $\omega_{0},\dots,\omega_{n-2}$. Then the following orbifold Bogomolov-Gieseker inequality holds:
@@
As another important application, we prove a quasi‑uniformization theorem for projective klt varieties with big canonical divisor. The Miyaoka-Yau inequality, originating in [82, 83], is a cornerstone in the geometry of surfaces of general type; see the introductions of [37, 43] for an overview. For a minimal projective klt variety $X$ of general type, the authors of [37] established the orbifold Miyaoka-Yau inequality
@@
Moreover, the equality case of (1.7) yields a quasi‑uniformization description [39]: the canonical model of $X$ is a singular quotient of the unit ball. Making use of the framework developed in this paper, we now extend this picture to varieties whose canonical divisor is big, but not necessarily nef.
@@
If equality in (1.8) holds, then the canonical model $X_{\mathrm{can}}$ of $X$ is a singular quotient of a projective manifold $Y$ that is covered by the unit ball $\mathbb{B}^{n}$.
The inequality (1.8) was established in the recent work [54]–a result that is new even in the smooth case. Here $\langle K_{X}^{n-2}\rangle$ denotes the non-pluripolar product; the precise definition of the intersection number appearing in (1.8) is given in [54, Section 4.1]. Together with the very recent preprint [59]—which proves that for a $K$-stable projective klt variety with big anti-canonical divisor, the Miyaoka-Yau type equality forces its anti-canonical model to be a singular quotient of $\mathbb{P}^{n}$—these results confirm that the inequality (1.8) developed in [54] provides an effective criterion for projective klt varieties with big canonical or anti-canonical divisors, and completes the picture
@@ 1.3. Overview and main technical components
Let $X$ be a compact Kähler klt space and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ a reflexive Higgs sheaf over $X_{\mathrm{reg}}$.
@@
    $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})\in\mathrm{Higgs}_{X_{\mathrm{reg}}}$ (resp. $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$) if and only if for every maximally quasi-étale cover $\gamma:Y\rightarrow X$, $\mathcal{E}_{Y}:=\gamma^{[*]}\mathcal{E}_{X}$ is locally free and $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ extends to a locally free Higgs sheaf $(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{Higgs}_{Y}$ (resp. $\mathrm{pHiggs}_{Y}$). Moreover, $(\mathcal{E}_{Y},\theta_{Y})$ is an extension of locally free Higgs sheaves belonging to $\mathrm{pHiggs}_{Y}$.
@@
    If $\mathcal{E}_{X}$ is locally free, then $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$ (resp. $\mathrm{pHiggs}_{X}$) if and only if for every resolution of singularities $\pi:\widehat{X}\rightarrow X$, the pullback $\pi^{*}(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{\widehat{X}}$ (resp. $\mathrm{pHiggs}_{\widehat{X}}$).
The main task is to prove that for any Jordan-Hölder filtration $\{\mathcal{E}_{i}\}_{i=0}^{l}$ of $\gamma^{[*]}(\mathcal{E}_{Y_{\mathrm{reg}}},\theta_{Y_{\mathrm{reg}}})$, each quotient sheaf $\mathcal{E}_{i}/\mathcal{E}_{i-1}$ is locally free and has vanishing Chern classes. Following the spirits of [38, Proposition 7.10], [49, Theorem 1.2] and [56, Proposition 3.10], the problem in our setting reduces to proving the following statement: if $(\mathcal{Q}_{Y_{\mathrm{reg}}},\theta_{\mathcal{Q}_{Y_{\mathrm{reg}}}})\in\operatorname{pHiggs}_{Y_{\mathrm{reg}}}$ is stable, then its reflexive extension $\mathcal{Q}_{Y}$ is locally free and all its Chern classes vanish. In the projective case [38, Proposition 7.10], this is handled by cutting down with hypersurface sections to reduce to surfaces, where reflexive sheaves are automatically locally free—a technique that is not available for general Kähler spaces. The central difficulty in the Kähler setting is therefore precisely to establish the local freeness of $\mathcal{Q}_{Y}$ in higher dimensions.
In Sections 3 and 4, we concentrate on proving the following characterization of polystable reflexive Higgs sheaves satisfying (1.3). This result not only yields the desired local freeness of $\mathcal{Q}_{Y}$ but also constructs the correspondence $\mu_{X_{\mathrm{reg}}}:\mathrm{pHiggs}_{X_{\mathrm{reg}}}\to\mathrm{sLSys}_{X_{\mathrm{reg}}}$.
@@
    A flat bundle $(E_{X_{\mathrm{reg}}},D_{X_{\mathrm{reg}}})$ is semisimple if and only if it admits a pluri-harmonic metric $H_{X_{\mathrm{reg}}}$, which is unique up to a positive constant on each direct summand.
@@
    A reflexive Higgs sheaf $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ belongs to $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$ if and only if $\mathcal{E}_{X_{\mathrm{reg}}}$ is locally free and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ arises from a harmonic bundle $(E_{X_{\mathrm{reg}}},D_{X_{\mathrm{reg}}},H_{X_{\mathrm{reg}}})$; here $H_{X_{\mathrm{reg}}}$ is unique up to obvious ambiguity. This is also equivalent to any one of the following three statements:
@@
    In Section 3, we prove the existence of a pluriharmonic metric for reflexive Higgs sheaves. This result was previously established in [28] for the case where $\theta_{X_{\mathrm{reg}}}$ is trivial (see also [41, 42]). Our proof adapts the Higgs version [74] of Bando-Siu’s argument [3] to study the Hermitian-Yang-Mills flow $\{H_{\mathrm{orb}}(t)\}$ on an orbifold modification constructed in [68]. Combining this with (1.3), we conclude that the limiting Hermitian-Einstein metric is pluri-harmonic. The overall strategy are presented in Section 3.1.
@@
    In the first part of Section 4.2, we prove part (1) of the theorem. The existence of pluri-harmonic metrics on semisimple flat bundles over the regular locus follows from the work of Corlette and Donaldson [17, 25] via maximally quasi‑étale covers; the remaining properties are obtained through the maximum principle, which is well adapted to our setting thanks to the extension theorem of Grauert and Remmert [32].
@@
    From Step 1 and Theorem 1.13 (1), we conclude that $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ admits a pluri-harmonic metric induced by a semisimple local system $E_{Y_{\mathrm{reg}}}$. Since $Y$ is maximally quasi-étale, $E_{Y_{\mathrm{reg}}}$ extends to a semisimple local system $E_{Y}$. To verify that $\gamma^{[*]}(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$ extends to a locally free Higgs sheaf on $Y$, it suffices to establish a descent result for Higgs bundles arising from harmonic bundles (Proposition 4.11), which we prove using the period map constructed in [19]. Details are given in Sections 4.2 and 4.3.